OpenStudy (jokersmart):

HELP, QUICKLY PLEASE. 5/x + 3/2x

OpenStudy (jokersmart):

\[\frac{ 5 }{ x }+\frac{ 3 }{ 2x }\]

OpenStudy (anonymous):

\[\frac{ 13}{ 2x}\]

OpenStudy (jokersmart):

Any explanation? It would be very helpful! @R5lover1997

OpenStudy (anonymous):

Ok first you have to multiply 5*2

OpenStudy (anonymous):

Which is 10, after that you had the 3 from the numerator